# Massively Parallel Minimum Spanning Tree in General Metric Spaces

### v1

@@ 4.2 Approximation Ratio
The edges of $E_{t}^{1}$ found in step 1 of Algorithm 1 by the modified Borůvka’s algorithm are all in $\hat{E}_{t}$. Take any set of vertices $A\in\widehat{\mathcal{P}}_{t}$. Both $E_{t}[A]$ and $\hat{E}_{t}[A]$ are spanning trees of $H[A]$. Observe that for an edge to be added by the modified Borůvka’s algorithm, it must have at some point been the minimum edge $e_{C}$ between a component $C$ and $H[A]\setminus C$. Assume for the sake of contradiction that $e_{C}$ is not in $\hat{E}_{t}$. In that case, we can replace an edge in $\hat{E}_{t}$ with $e_{C}$ to create a spanning tree respecting $(G_{\leq\alpha t}\mid\widehat{\mathcal{P}}_{t})$ that is smaller than $\hat{T}$, which is a contradiction. More precisely, as $\hat{E}_{t}$ is a spanning tree on $H[A]$, it contains a path $P$ from one endpoint of $e_{C}$ to the other. Therefore, $P$ must cross between $C$ and $H[A]\setminus C$ using some edge $e^{\prime}$. Then, $\hat{T}\setminus\{e_{C}\}\cup\{e^{\prime}\}$ is a smaller spanning tree w.r.t. $(G_{\leq\alpha t}\mid\widehat{\mathcal{P}}_{t})$, which contradicts the fact $\hat{T}$ is the MST.
Now, we analyze the edges $E_{t}^{2}$ found in step 1, i.e. the edges that join the remaining components inside a set of $\widehat{\mathcal{P}}_{t}$ that were not joined by step 1. Let $N^{(r^{\prime})}_{A}$ be the number of surviving components after $r^{\prime}$ rounds of the modified Borůvka algorithm, where $0\leq r^{\prime}\leq r$. We say a component has survived a round of the modified Borůvka algorithm if it does not join any other component. $N^{(0)}_{A}$ is simply the number of vertices in $H[A]$, and $N^{(r)}_{A}$ components survive in $A$ in the end. Using the same analysis as the one for the number of surviving connected components of leader compression (see the proof of Claim 4.6), we can show that
@@
$=2^{-\Omega(r)}\alpha t\cdot\left(\lvert\widehat{\mathcal{P}}_{t/\alpha}\rvert-\lvert\widehat{\mathcal{P}}_{t}\rvert\right),$
@@
$\leq w(\hat{T})+2^{-\Omega(r)}\sum_{t}\alpha t\cdot\left(\lvert\widehat{\mathcal{P}}_{t/\alpha}\rvert-\lvert\widehat{\mathcal{P}}_{t}\rvert\right)$
@@
where the next to last inequality is by Claim 4.6, and the last inequality by Claim 4.7 and the choice of $r$. This concludes the proof. ∎
@@ 4.3 Runtime Analysis and Implementation Details
Algorithm 1 can be implemented in $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of the MPC model, with $O(n^{\delta})$ space per machine and $\widetilde{O}(n^{2})$ total space, where $10<\delta<$ 1.
@@
    Sorting [goodrich2011sorting]: A sequence of $N$ elements can be sorted in $O\left(\frac{1}{\delta}\right)$ rounds using $O(N^{\delta})$ space per machine and $O(N)$ total space.
@@
[…]
@@
The tree $T$ in Algorithm 1 can be computed from $\hat{\mathcal{P}}$ in $O\left(\frac{1}{\delta}\left(\log\frac{1}{\varepsilon}+\log\log n\right)\right)$ rounds in the MPC model.
@@ 5 A Lower Bound for Metric MST
In this section, we prove Theorem 2.
@@
Throughout the proof, we assume that $\frac{1}{\varepsilon}=O(n^{\delta})$, where $O(n^{\delta})$ is the local space on each machine and $10<\delta<$ 1. This is without loss of generality since any $(1+\frac{1}{n})$-approximation is a $(1+\frac{1}{n^{\delta}})$-approximation, and $\log n=\Theta(\log n^{\delta})$.
To make the reduction, we present an algorithm that given a cycle of length $n$, with high probability outputs the order in which the vertices appear in the cycle. Given any input graph, the output of the algorithm can be used to validate if the input is a cycle of length $n$. That is, given the output ordering $u_{0},\ldots u_{n-1}$, one can simply check whether there are $n$ edges, one between $u_{i}$ and $u_{(i+1)\bmod n}$ for each $10\leq i\leq n-$ 1.
@@
Observe, the above implies that for each vertex $u$, we can find the subpath of length $\frac{2}{\varepsilon}$ that is centered at $u$ by gathering all the vertices in its $\frac{1}{\varepsilon}$-neighborhood in one machine. A cycle of length $L$ can then be broken down into $\frac{1}{\varepsilon}$ cycles of length $\varepsilon L$ by removing all the edges, and connecting each vertex to the two vertices that were at distance exactly $\frac{1}{\varepsilon}$ (Figure 2). More formally,
@@
Given a graph $G$, consisting of $\frac{1}{\varepsilon^{k}}$ cycles $C_{0},\ldots,C_{1/\varepsilon^{k}-1}$ of length $\varepsilon^{k}n$, and a list of vertices $u^{0}_{0},\ldots,u^{1/\varepsilon^{k}-1}_{0}$, such that the cycle $C_{i}$ contains vertices $u^{i}_{0},\ldots,u^{i}_{\varepsilon^{k}n-1}$ (in that order), using $O(n^{\delta})$ space per machine, $\textnormal{poly}(n)$ total space, and $O(R)$ rounds, we can construct a new graph $G^{\prime}$ on the same vertex set, such that for each $i$, there are $\frac{1}{\varepsilon}$ cycles $C^{i}_{0},\ldots,C^{i}_{1/\varepsilon-1}$ in $G^{\prime}$, such that $C^{i}_{j}$ consists of vertices $u^{i}_{j},u^{i}_{j+1/\varepsilon},u^{i}_{j+2/\varepsilon},\ldots,u^{i}_{j+(\varepsilon^{k+1}n-1)/\varepsilon}$ in that order. Furthermore, we can compute the list of vertices $u^{i}_{j}$, where $10\leq i\leq 1/\varepsilon^{k}-$ 1 and $10\leq j\leq 1/\varepsilon-$ 1. That is, one vertex $u^{i}_{j}$ per cycle $C^{i}_{j}$, such that for each $i$ they form a subpath in $C_{i}$.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$$v_{11}$$v_{12}$$v_{13}$$v_{14}$$v_{15}$(a)$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$$v_{11}$$v_{12}$$v_{13}$$v_{14}$$v_{15}$$v_{1}$$v_{4}$$v_{7}$$v_{10}$$v_{13}$$v_{2}$$v_{5}$$v_{8}$$v_{11}$$v_{14}$$v_{3}$$v_{6}$$v_{9}$$v_{12}$$v_{15}$(b)
@@
First, we show how to utilize Claim 5.3 to find the order of the cycles. Assume for simplicity that $\frac{1}{\varepsilon}$ is an integer, this can be done by taking a smaller parameter $\varepsilon^{\prime}$ such that $\frac{\varepsilon}{2}\leq\varepsilon^{\prime}\leq\varepsilon$. Also, we can assume that the number of vertices $n$ is a power of $\frac{1}{\varepsilon}$. This is without loss of generality since we can round up $n$ to the nearest power of $\frac{1}{\varepsilon}$, say $n^{\prime}$, by adding $n^{\prime}-n$ extra edges anywhere in the cycle. As a result, the number of vertices grows by at most a factor of $\frac{1}{\varepsilon}$, and is still $\textnormal{poly}(n)$.
Given a cycle of length $n=\frac{1}{\varepsilon^{N}}$, we can find the order of the vertices by recursively applying Claim 5.3. Take any vertex $u_{0}$, and let $u_{0},\ldots,u_{n-1}$ be the order of the cycle. To start, we can feed the graph and $u_{0}$ to Claim 5.3. We get a new graph consisting of the cycles $C_{0},\ldots,C_{1/\varepsilon-1}$ of length $\frac{1}{\varepsilon^{N-1}}$. Note that we do not know what the cycles are. Rather, we know of only one vertex in each cycle $C_{i}$, namely $u_{i}$, where $u_{0},\ldots,u_{1/\varepsilon-1}$ form a subpath in the cycle.
After computing the order of vertices in each cycle $C_{i}$ recursively, we can recover the order in the original cycle by “shuffling” the orders of $C_{i}$’s as follows. First, we add the first vertex in each cycle to the order, i.e. $u_{0},\ldots,u_{1/\varepsilon-1}$. Then we add the second vertex of each cycle to the order, these are the $\frac{1}{\varepsilon}$ vertices that follow $u_{0},\ldots,u_{1/\varepsilon-1}$. Then, the third, and so on (Figure 2).
@@
The remainder of this section is devoted to proving Claims 5.2 and 5.3. We begin by presenting some auxiliary tools and claims used in the proofs.
@@
Proof by picture: see Figures 3 and 4.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$$v_{11}$$v_{12}$$v_{13}$$v_{14}$$v_{15}$$v_{16}$$v_{17}$$v_{18}$$v_{19}$$v_{20}$(a)$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$$v_{11}$$v_{12}$$v_{13}$$v_{14}$$v_{15}$$v_{16}$$v_{17}$$v_{18}$$v_{19}$$v_{20}$(b)
Figure 3: A possible outcome of a two-switch. (b) is obtained from (a) by doing a two-switch on $(v_{1},v_{2})$ and $(v_{16},v_{15})$. Note that $v_{1}$ and $v_{16}$ are on different cycles here (case 1 in Claim 5.3).
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$(a)$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$(b)$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$(c)
Figure 4: Two possible outcomes of a two-switch. (b) is obtained from (a) by doing a two-switch on $(v_{1},v_{2})$ and $(v_{5},v_{6})$. Note that $v_{6}$ and $v_{2}$ are on different paths from $v_{1}$ to $v_{5}$ (case 2 in Claim 5.3). (c) is obtained from (a) by doing a two-switch on $(v_{1},v_{2})$ and $(v_{6},v_{5})$. Observe that $v_{5}$ and $v_{2}$, are on the same path from $v_{1}$ to $v_{6}$ (case 3 in Claim 5.3).
Given a graph $G$ with $n$ vertices that is a disjoint union of $c$ cycles, we relate it to MST approximation in $(1,2)$-metrics as follows. Let two points in the metric have distance $1$ if there is an edge between them in $G$, and distance $2$ otherwise. An MST of this metric has weight $n+c-2$, as it uses $n-c$ edges of weight $1$, and $c-1$ edges of weight $2$. Alternatively, this can be viewed as turning $G$ into a forest by removing $c$ edges, and then connecting the components to build the MST. Similarly, for an approximate MST, the problem can be viewed as turning $G$ into a forest by removing not many more than $c$ edges, and then connecting the components using edges of weight at most $2$.
@@
Observe that any spanning tree that removes $k$ edges of $G$, contains $n-k$ weight-$1$ edges, and $k-1$ weight-$2$ edges. Therefore it has weight $n+k-2$ edges. As mentioned, the MST has weight $n+c-2$ edges since it removes $c$ weight-$1$ edges, one from each cycle. Hence, any $(1+\varepsilon)$-approximate MST has weight at most
@@
and removes at most $c+\varepsilon(n+c-2)$ weight-$1$ edges. ∎
Given an MPC algorithm that finds a $(1+\varepsilon)$-approximation of the MST in $(1,2)$-metric, we examine the probability of each weight-$1$ edge getting removed. This probability is used to determine the $\frac{1}{\varepsilon}$-neighborhood of a vertex in Claim 5.2. We keep referring to the subgraph of weight-$1$ edges as $G$. We show that by a random reordering of the vertex IDs, any algorithm can be transformed such that for any two isomorphic edges, the probability of being removed by the algorithm is the same.
@@
Given a graph $G$, two edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ are isomorphic, if there is an automorphism of $G$ that maps one to the other. That is, if there is a bijection $\phi:V(G)\to V(G)$, such that $\phi(u_{1})=u_{2}$, $\phi(v_{1})=v_{2}$, and for any pair of vertices $u$ and $v$, there is an edge $(u,v)$ if and only if there an edge $(\phi(u),\phi(v))$.
@@
[…]
@@
With that we are ready to prove Claim 5.2.
Proof of Claim 5.2.
Given $\frac{1}{\varepsilon^{k}}$ cycles of length $\varepsilon^{k}n$, the goal is to find the $\frac{1}{\varepsilon}$-neighborhood of each vertex. We are given an algorithm $\mathcal{A}$ that computes a $(1+\varepsilon)$-approximate MST for $(1,2)$-metrics with high probability, using $O(n^{\delta})$ space per machine and $\textnormal{poly}(n)$ total space, in $R$ rounds. Due to Lemma 5.9, we can assume that $\mathcal{A}$ removes isomorphic edges with the same probability. Specifically, in a graph that is a disjoint union of cycles, the length of a cycle determines the removal probability of its edges.
@@
The goal is to distinguish the case where $u_{1}$ and $u_{2}$ are at most $D$ edges apart, from the other cases. Roughly, if $u_{1}$ and $u_{2}$ are close, then doing the two-switch with the appropriate neighbors creates a small cycle. The edges of a small cycle have a larger removal probability than an average edge, and hence can be identified. More formally, we examine the removal probabilities of all the edges after the two-switch, and use them to assert if $u_{1}$ and $u_{2}$ have distance at most $D$. A summary of the different scenarios appears in Table 1.
@@
The edges in $E_{1}$ are all isomorphic, and there are $n-\varepsilon^{k}n\geq\frac{n}{2}$ of them. $\mathcal{A}$ computes a $(1+\varepsilon)$-approximation of the MST. Therefore, by Claim 5.7, with high probability the number of edges it removes is at most (here, $c$ is the number of cycles in $G^{\prime}$)
@@
and the other two with $>\frac{1}{20\varepsilon}$ edges. (4-c)
@@
[…]
@@
Now, we group the edges based on their approximated removal probabilities. We call two edges $e_{1}$ and $e_{2}$ close if $\lvert\hat{p}_{e_{1}}-\hat{p}_{e_{2}}\rvert<\frac{\varepsilon}{2}$. We define the groups as the transitive closure of closeness. That is, we let $e$ and $e^{\prime}$ be in the same group if there is a chain of edges $e_{0},e_{1},\ldots,e_{l}$, such that $e_{0}=e$, $e_{l}=e^{\prime}$, and $\lvert\hat{p}_{e_{i}}-\hat{p}_{e_{i+1}}\rvert\leq\frac{\varepsilon}{2}$ for $10\leq i\leq l-$ 1. These groups can be computed in $O(1)$ rounds of MPC, by sorting the edges based on their approximated removal probabilities.
@@
A pair of vertices $u_{1}$ and $u_{2}$ are within distance $\frac{1}{20\varepsilon}$ of each other if and only if for some choice of neighbors, $v_{1}$ and $v_{2}$, case 3 happens. Given the characterizations of the groups, and the different cases (Table 1), we can tell when $(u_{1},v_{1})$ and $(u_{2},v_{2})$ are in case 3 (i.e. 3-a, 3-b, or 3-c) as follows.

### v2

@@ 4.2 Approximation Ratio
[…]
Now, we analyze the edges $E_{t}^{2}$ found in step 1, i.e. the edges that join the remaining components inside a set of $\widehat{\mathcal{P}}_{t}$ that were not joined by step 1. Let $N^{(r^{\prime})}_{A}$ be the number of surviving components after $r^{\prime}$ rounds of the modified Borůvka algorithm, where $0\leq r^{\prime}\leq r$. We say a component has survived a round of the modified Borůvka algorithm if it does not join any other component. $N^{(0)}_{A}$ is simply the number of vertices in $H[A]$, and $N^{(r)}_{A}$ components survive in $A$ in the end. Using the same analysis as the one for the number of surviving connected components of leader compression (see the proof of Claim˜4.6), we can show that
@@
[…]
@@
[…]
@@
where the next to last inequality is by Claim˜4.6, and the last inequality by Claim˜4.7 and the choice of $r$. This concludes the proof. ∎
@@ 4.3 Runtime Analysis and Implementation Details
Algorithm˜1 can be implemented in $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of the MPC model, with $O(n^{\delta})$ space per machine and $\widetilde{O}(n^{2})$ total space, where $0<\delta<1$.
@@
    Sorting [GSZ11]: A sequence of $N$ elements can be sorted in $O\left(\frac{1}{\delta}\right)$ rounds using $O(N^{\delta})$ space per machine and $O(N)$ total space.
@@
First, we need to construct $\mathcal{P}^{0}$. The data is replicated $L$ times, one copy per level. Then, for each level $t=\alpha^{k}$, the partition $\mathcal{P}^{0}_{t}$ is built independently. A set of $O(n^{1-\delta})$ machines $M_{u}$ are created for each vertex $u$. Using sorting the edges $(u,v)$ for all other $v$ are transported to $M_{u}$. Then, a predesignated machine in $M_{u}$ draws the delay $\delta_{u}$ and communicates it to the rest of $M_{u}$. At this point, $M_{u}$ creates a tuple $\big(v,\ u,\ w(u,v)-\delta_{u}\big)$ for each vertex $v$. This tuple represents $(\textnormal{vertex},\ \textnormal{potential center},\ \textnormal{shifted distance})$. Finally, using sorting, $M_{u}$ collects all the tuples $\big(u,\ v,\ w(u,v)-\delta_{v}\big)$ and identifies the vertex $v$ with the smallest shifted distance as its center. The component IDs in $\mathcal{P}^{0}_{t}$ are simply the vertex ID of the center for that component.
@@
The tree $T$ in Algorithm˜1 can be computed from $\hat{\mathcal{P}}$ in $O\left(\frac{1}{\delta}\left(\log\frac{1}{\varepsilon}+\log\log n\right)\right)$ rounds in the MPC model.
@@ 5 A Lower Bound for Metric MST
In this section, we prove Theorem˜2, restated below. See 2
@@
Throughout the proof, we assume that $\frac{1}{\varepsilon}=O(n^{\delta})$, where $O(n^{\delta})$ is the local space on each machine and $0<\delta<1$. This is without loss of generality since any $(1+\frac{1}{n})$-approximation is a $(1+\frac{1}{n^{\delta}})$-approximation, and $\log n=\Theta(\log n^{\delta})$.
To make the reduction, we present an algorithm that given a cycle of length $n$, with high probability outputs the order in which the vertices appear in the cycle. Given any input graph, the output of the algorithm can be used to validate if the input is a cycle of length $n$. That is, given the output ordering $u_{0},\ldots u_{n-1}$, one can simply check whether there are $n$ edges, one between $u_{i}$ and $u_{(i+1)\bmod n}$ for each $0\leq i\leq n-1$.
@@
Observe, the above implies that for each vertex $u$, we can find the subpath of length $\frac{2}{\varepsilon}$ that is centered at $u$ by gathering all the vertices in its $\frac{1}{\varepsilon}$-neighborhood in one machine. A cycle of length $L$ can then be broken down into $\frac{1}{\varepsilon}$ cycles of length $\varepsilon L$ by removing all the edges, and connecting each vertex to the two vertices that were at distance exactly $\frac{1}{\varepsilon}$ (Figure˜2). More formally,
@@
Given a graph $G$, consisting of $\frac{1}{\varepsilon^{k}}$ cycles $C_{0},\ldots,C_{1/\varepsilon^{k}-1}$ of length $\varepsilon^{k}n$, and a list of vertices $u^{0}_{0},\ldots,u^{1/\varepsilon^{k}-1}_{0}$, such that the cycle $C_{i}$ contains vertices $u^{i}_{0},\ldots,u^{i}_{\varepsilon^{k}n-1}$ (in that order), using $O(n^{\delta})$ space per machine, $\textnormal{poly}(n)$ total space, and $O(R)$ rounds, we can construct a new graph $G^{\prime}$ on the same vertex set, such that for each $i$, there are $\frac{1}{\varepsilon}$ cycles $C^{i}_{0},\ldots,C^{i}_{1/\varepsilon-1}$ in $G^{\prime}$, such that $C^{i}_{j}$ consists of vertices $u^{i}_{j},u^{i}_{j+1/\varepsilon},u^{i}_{j+2/\varepsilon},\ldots,u^{i}_{j+(\varepsilon^{k+1}n-1)/\varepsilon}$ in that order. Furthermore, we can compute the list of vertices $u^{i}_{j}$, where $0\leq i\leq 1/\varepsilon^{k}-1$ and $0\leq j\leq 1/\varepsilon-1$. That is, one vertex $u^{i}_{j}$ per cycle $C^{i}_{j}$, such that for each $i$ they form a subpath in $C_{i}$.
[…]
@@
First, we show how to utilize Claim˜5.3 to find the order of the cycles. Assume for simplicity that $\frac{1}{\varepsilon}$ is an integer. This can be done by taking a smaller parameter $\varepsilon^{\prime}$ such that $\frac{\varepsilon}{2}\leq\varepsilon^{\prime}\leq\varepsilon$. Also, we can assume that the number of vertices $n$ is a power of $\frac{1}{\varepsilon}$. This is without loss of generality since we can round up $n$ to the nearest power of $\frac{1}{\varepsilon}$, say $n^{\prime}$, by adding $n^{\prime}-n$ extra edges anywhere in the cycle. As a result, the number of vertices grows by at most a factor of $\frac{1}{\varepsilon}$, and is still $\textnormal{poly}(n)$.
Given a cycle of length $n=\frac{1}{\varepsilon^{N}}$, we can find the order of the vertices by recursively applying Claim˜5.3. Take any vertex $u_{0}$, and let $u_{0},\ldots,u_{n-1}$ be the order of the cycle. To start, we can feed the graph and $u_{0}$ to Claim˜5.3. We get a new graph consisting of the cycles $C_{0},\ldots,C_{1/\varepsilon-1}$ of length $\frac{1}{\varepsilon^{N-1}}$. Note that we do not know what the cycles are. Rather, we know of only one vertex in each cycle $C_{i}$, namely $u_{i}$, where $u_{0},\ldots,u_{1/\varepsilon-1}$ form a subpath in the cycle.
After computing the order of vertices in each cycle $C_{i}$ recursively, we can recover the order in the original cycle by “shuffling” the orders of $C_{i}$’s as follows. First, we add the first vertex in each cycle to the order, i.e. $u_{0},\ldots,u_{1/\varepsilon-1}$. Then we add the second vertex of each cycle to the order, these are the $\frac{1}{\varepsilon}$ vertices that follow $u_{0},\ldots,u_{1/\varepsilon-1}$. Then, the third, and so on (Figure˜2).
@@
The remainder of this section is devoted to proving Claims˜5.2 and 5.3. We begin by presenting some auxiliary tools and claims used in the proofs.
@@
Proof by picture: see Figures˜3 and 4.
[…]
Figure 3: A possible outcome of a two-switch. (b) is obtained from (a) by doing a two-switch on $(v_{1},v_{2})$ and $(v_{16},v_{15})$. Note that $v_{1}$ and $v_{16}$ are on different cycles here (case 1 in Claim˜5.3).
[…]
Figure 4: Two possible outcomes of a two-switch. (b) is obtained from (a) by doing a two-switch on $(v_{1},v_{2})$ and $(v_{5},v_{6})$. Note that $v_{6}$ and $v_{2}$ are on different paths from $v_{1}$ to $v_{5}$ (case 2 in Claim˜5.3). (c) is obtained from (a) by doing a two-switch on $(v_{1},v_{2})$ and $(v_{6},v_{5})$. Observe that $v_{5}$ and $v_{2}$, are on the same path from $v_{1}$ to $v_{6}$ (case 3 in Claim˜5.3).
Given a graph $G$ with $n$ vertices that is a disjoint union of $c$ cycles, we relate it to MST approximation in $(1,2)$-metrics as follows. Let two points in the metric have distance $1$ if there is an edge between them in $G$, and distance $2$ otherwise. An MST of this metric has weight n+c2n+c-2, as it uses ncn-c edges of weight $1$, and c1c-1 edges of weight $2$. Alternatively, this can be viewed as turning $G$ into a forest by removing $c$ edges, and then connecting the components to build the MST. Similarly, for an approximate MST, the problem can be viewed as turning $G$ into a forest by removing not many more than $c$ edges, and then connecting the components using edges of weight at most $2$.
@@
Observe that any spanning tree that removes $k$ edges of $G$, contains nkn-k weight-11 edges, and k1k-1 weight-22 edges. Therefore, it has weight n+k2n+k-2 edges. As mentioned, the MST has weight n+c2n+c-2 edges since it removes $c$ weight-11 edges, one from each cycle. Hence, any $(1+\varepsilon)$-approximate MST has weight at most
@@
and removes at most $c+\varepsilon(n+c-2)$ weight-11 edges. ∎
Given an MPC algorithm that finds a $(1+\varepsilon)$-approximation of the MST in $(1,2)$-metric, we examine the probability of each weight-$1$ edge getting removed. This probability is used to determine the $\frac{1}{\varepsilon}$-neighborhood of a vertex in Claim˜5.2. We keep referring to the subgraph of weight-$1$ edges as $G$. We show that by a random reordering of the vertex IDs, any algorithm can be transformed such that for any two isomorphic edges, the probability of being removed by the algorithm is the same.
@@
Given a graph $G$, two edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ are isomorphic if there is an automorphism of $G$ that maps one to the other. That is, if there is a bijection $\phi:V(G)\to V(G)$, such that $\phi(u_{1})=u_{2}$, $\phi(v_{1})=v_{2}$, and for any pair of vertices $u$ and $v$, there is an edge $(u,v)$ if and only if there an edge $(\phi(u),\phi(v))$.
@@
$=\operatorname*{\textnormal{Pr}}(\mathcal{A}^{\prime}\textnormal{ removes }e^{\prime}).$
@@
With that, we are ready to prove Claim˜5.2.
Proof of Claim˜5.2.
Given $\frac{1}{\varepsilon^{k}}$ cycles of length $\varepsilon^{k}n$, the goal is to find the $\frac{1}{\varepsilon}$-neighborhood of each vertex. We are given an algorithm $\mathcal{A}$ that computes a $(1+\varepsilon)$-approximate MST for $(1,2)$-metrics with high probability, using $O(n^{\delta})$ space per machine and $\textnormal{poly}(n)$ total space, in $R$ rounds. Due to Lemma˜5.9, we can assume that $\mathcal{A}$ removes isomorphic edges with the same probability. Specifically, in a graph that is a disjoint union of cycles, the length of a cycle determines the removal probability of its edges.
@@
The goal is to distinguish the case where $u_{1}$ and $u_{2}$ are at most $D$ edges apart, from the other cases. Roughly, if $u_{1}$ and $u_{2}$ are close, then doing the two-switch with the appropriate neighbors creates a small cycle. The edges of a small cycle have a larger removal probability than an average edge, and hence can be identified. More formally, we examine the removal probabilities of all the edges after the two-switch, and use them to assert if $u_{1}$ and $u_{2}$ have distance at most $D$. A summary of the different scenarios appears in Table˜1.
@@
The edges in $E_{1}$ are all isomorphic, and there are $n-\varepsilon^{k}n\geq\frac{n}{2}$ of them. $\mathcal{A}$ computes a $(1+\varepsilon)$-approximation of the MST. Therefore, by Claim˜5.7, with high probability the number of edges it removes is at most (here, $c$ is the number of cycles in $G^{\prime}$)
@@
and the other two with $\frac{1}{20\varepsilon}$ edges. (4-c)
@@
$\operatorname*{\textnormal{Pr}}(\lvert\hat{p}_{e}-p_{e}\rvert>\varepsilon/4)\leq 2e^{-32S\varepsilon^{2}}\leq n^{-c},$ (7)
@@
Now, we group the edges based on their approximated removal probabilities. We call two edges $e_{1}$ and $e_{2}$ close if $\lvert\hat{p}_{e_{1}}-\hat{p}_{e_{2}}\rvert<\frac{\varepsilon}{2}$. We define the groups as the transitive closure of closeness. That is, we let $e$ and $e^{\prime}$ be in the same group if there is a chain of edges $e_{0},e_{1},\ldots,e_{l}$, such that $e_{0}=e$, $e_{l}=e^{\prime}$, and $\lvert\hat{p}_{e_{i}}-\hat{p}_{e_{i+1}}\rvert\leq\frac{\varepsilon}{2}$ for $0\leq i\leq l-1$. These groups can be computed in $O(1)$ rounds of MPC, by sorting the edges based on their approximated removal probabilities.
@@
A pair of vertices $u_{1}$ and $u_{2}$ are within distance $\frac{1}{20\varepsilon}$ of each other if and only if for some choice of neighbors, $v_{1}$ and $v_{2}$, case 3 happens. Given the characterizations of the groups, and the different cases (Table˜1), we can tell when $(u_{1},v_{1})$ and $(u_{2},v_{2})$ are in case 3 (i.e. 3-a, 3-b, or 3-c) as follows.